# SPICE: A Synergistic, Precise, Iterative, and Customizable Image Editing Workflow

### v1

@@ 3.1 Benchmark Results
Following Sheynin et al. (2024), we display CLIPdir and CLIPout in Table 1 to compare CLIP text-image direction similarity and CLIP output similarity. We can see that our workflow achieves the highest scores on the EditEval benchmark. In addition, we compare the edited images by baseline methods and our workflow on all six different editing categories for visual comparison. Figure 3 shows the failure patterns of baseline methods including undesired global color shift of the grass (Object Addition), removing the lens flare outside the masked region (Object Replacement), inability to remove multiple humans from the image (Object Removal), and wrong anatomy of the polar bear (Action Change). In contrast, our workflow maintains global color, keeps details outside the mask untouched, removes objects as requested, and generates animals with correct anatomy.
Table 1: Our workflow achieves top performance in two quantitative metrics on the EditEval benchmark. We show mean $\pm$ standard deviation calculated across all images.
@@
To provide a comprehensive qualitative evaluation, we further conduct a human evaluation to compare results from different models. We ask a total of 9 annotators to select their preferred image from an image pair, where one image comes from our workflow and another image comes from one of the baseline methods. To avoid introducing any bias as the workflow designers, we simply ask the annotators to select the image that better follows the editing instruction (more details in Appendix B).
Figure 4 shows the human evaluation results. Our workflow is predominantly preferred. Although at a much lower chance than baseline methods, our workflow does sometimes fail to follow instructions, as indicated by the “both bad” cases. In Section 3.3, we show how the flexibility of our workflow allows us to overcome this limitation easily.
Figure 4: Our workflow is preferred by the annotators over any baseline method. In this evaluation, we use only a single editing step and fixed hyperparameters for our workflow, and thus our workflow can still fail (“both bad” cases). In Section 3.3, we show how relaxing these constraints can dramatically improve results.
@@
In Appendix C, we show that our workflow outperforms baselines on challenging examples from two other popular benchmarks (Emu Edit and MagicBrush). On the challenging examples, our workflow also outperforms SeedEdit (Shi et al., 2024a), a latest commercial baseline from Doubao AI.
@@
While FLUX.1 [dev] is a strong image generation backbone, we demonstrate that our workflow outperforms the backbone model FLUX.1 [dev] by an ablation study. Also, we designed our workflow such that it mitigates many issues of the specialized inpainting model FLUX.1 [dev] Fill [^11] https://huggingface.co/black-forest-labs/FLUX.1-Fill-dev. Hence, the better performance of our workflow than baseline methods does not result from simply selecting a stronger backbone. Results can be found in Appendix D.
@@
Despite the superior performance, our workflow imposes minimal burden on the users. While the users need to provide both the masks and the hints, the two inputs can be casually sketched, because our workflow is robust to missing details and inaccurate shapes. Examples of excellent editing results from simple user inputs are shown in Figure 5.
Original Hinted Silhouette Mask Result
@@
Figure 6: Our workflow delivers precise and repeatable edits, where the generated objects align with user specifications. The top images compare the user-specified color hints with the generated objects, and the bottom plots show the percentage errors. On each plot, we show the mean and standard deviation of percentage errors across 10 random seeds. All mean percentage errors are close to 0.
3.2 Precise Editing
@@
To quantify the precision of our workflow, we add objects while specifying 5 properties, including size, location, rotation, color, and aspect ratio (Figure 6). To specify the value of each property, we use a hint in the form of a color patch. In the edited image, we measure the property value of the added object and calculate the percentage error against the user-specified property value (Appendix E). We use 10 random seeds to account for generative variability.
@@
Figure 6 shows that the mean of percentage errors for each specified property value is close to 0, with a small standard deviation across random seeds. There are still systematic errors, such as the height of the apple being consistently 10% larger than specified. However, for the apple, the height is larger because the apple’s stem is unspecified by the hint. Overall, the results confirm that our workflow delivers precise, repeatable edits critical for real-world scenarios.
3.3 Iterative Editing
@@
In this subsection, we consider two iterative editing regimes, one using fewer than 5 editing steps, another more than 100. Multiple editing steps are often necessary, as one single editing or generation step usually fails to align with complicated user requirements (further discussed in Section 3.4).
@@
In the low-steps regime, if we are not satisfied with results from the first editing step, we are able to improve the result using another editing step with our workflow. In Section 3.1, we identify that our workflow does not work well for a few cases in single-step editing. In these cases, a large proportion of the edited region is satisfactory, and we want to fix the small remaining proportion. If we are using one of the baseline methods, we can only run the model multiple times with the same source image and editing prompt, hoping to get a better result with a different random seed (Appendix F). However, with our workflow, we can simply improve the result by recycling the output from the first step (Figure 7).
@@
In the high-steps regime, while many previous methods support multi-step editing, heavy artifacts are generated and propagated in each step, disqualifying these methods for long editing tasks. For example, when Emu Edit is prompted to edit an image multiple times without changing a red bench in the scene [^12] https://emu-edit.metademolab.com/assets/videos/dog.mp4, the image area containing the bench is still edited, and the quality of this area becomes worse with each edit. The edges of planks that constitute the bench are increasingly blurred and deformed. In contrast, our workflow steadily improves the image after multiple steps. In LABEL:subfig:iterative-heart, we show a result after 40 diverse editing steps, including outpainting, structural edits, upscaling, relighting, and composition adjustment (forming a heart shape using the sky, the crescent, the claw, and the branches). Interestingly, to relight characters according to the environment, we do not need to provide any color hints. We also do not need a specialized relighting model, such as IC-Light (Zhang et al., 2025). Full steps are shown in Appendix G. In LABEL:subfig:iterative-maid, we start with a model-generated image with various errors in shadows, objects, and anatomy. Using the same model as the backbone in our workflow, we fix these errors one by one across more than 100 editing steps. This image is produced by BoleroMix (Pony) v1.41, a checkpoint derived from SDXL. More examples of final edited results in different styles can be found in Appendix H, where usually 100 to 200 editing steps are performed to achieve the desired level of correctness in details. Each image is created from scratch within 4 hours in one sitting.
Original Step 1 Step 4
Refer to caption Refer to caption Refer to caption
@@
3.4 Customizable Editing
@@
Prompt-based image generation models frequently fail to fulfill their promise that users can create content according to their own will. Users are frustrated for two main reasons.
@@
First, prompts are hard to design and often cannot instruct the model to produce complex output. A user assumes that the model follows the prompt, but the model does not interpret the prompt as humans do. One example is that DALL·E 3 (Betker et al., 2023) in November 2024 still fails to generate a rabbit with four ears or a violin without a bridge [^13] https://cs.nyu.edu/~davise/papers/DALL-E-Parts/PartsNovember24/DALL-E-Parts-November24.html, as shown in LABEL:subfig:dalle. Another example is that DALL·E 3 fails to interpret relations or numbers, such as a potato under a spoon or five fish (Conwell et al., 2024).
@@
Secondly, the innate randomness of the model dramatically increases the generation cost or even forbids the generation of complicated images. Suppose a user wants to generate an image with 10 different objects at various locations on this image. If each object has a 50% chance to be generated perfectly, the success rate of all objects being perfect is below 0.1%. In real scenarios, the perfection rate for each object will only be lower, and the number of objects larger. Then, generating such an image is practically impossible.
With SPICE, instead of desperately engineering prompts or varying the random seed, users handle these difficulties by tuning 3 hyperparameters, namely context size, denoising strength, and Canny model steps. We first demonstrate the effect of each hyperparameter (LABEL:fig:hyperparameters). Then, we show how these hyperparameters can help the user reliably generate an image that DALL·E 3 cannot (LABEL:fig:rabbit).
Context Size.
During inpainting, a diffusion model generates content consistent with what is already in the context. In LABEL:subfig:context-size, the user requires another yellow vegetable to be added to the center of the image. The prompt is “yellow vegetables on an orange background”. The color hint is the yellow circle in the center. The model’s interpretation of the color hint depends on the selected context. When a small context includes none of the other vegetables, the model generates multiple vegetables from the single yellow circle. When a medium context includes corners of the other vegetables, the model generates one vegetable. When a large context includes all other vegetables, the model fails to generate a new vegetable. When the context includes one vegetable either on the top left or the top right corner, the model generates a vegetable with the same direction of shade. While we do not theoretically explain the failure when using large context sizes, we observe it frequently in other editing tasks. We also observe the failure to add objects using the “whole image” mode (Section 2.1), an observation that motivated our design of context dots.
Denoising Strength.
The model balances hinted colors and realistic shadows at a medium denoising strength. In LABEL:subfig:denoising-strength, the user requires a Rubik’s cube with certain face colors to be added to the image. Because the colors cannot be succinctly described by words, the user inserts a reference image as the hint. With a low denoising strength, the colors are preserved, but the shadows are not generated. We observe the opposite with a high denoising strength.
Canny Model Steps.
The model balances realistic details and hinted edges at a medium number of Canny model steps. In LABEL:subfig:canny-model-steps, the user requires a crescent to be added to the sky, but the prompt is simply “a moon”. With few Canny model steps, the moon looks realistic, but its shape is wrong. We observe the opposite with many Canny model steps.
By slightly tuning the 3 hyperparameters (Appendix I), we are able to generate LABEL:subfig:customizable, an image with a rabbit of four ears, holding a violin without a bridge with the left hand and a spoon with the right hand. There is a potato under the spoon and five fish in the air. In contrast, DALL·E 3 cannot generate even one of the components on this image.
4 Related Work
Image Editing Methods.
Diffusion model-based image editing techniques can be categorized based on the type of inputs used to guide image generation (Croitoru et al., 2023; Yang et al., 2023; Huang et al., 2024). Most existing methods rely primarily on textual prompts (Brooks et al., 2023; Fu et al., 2024), while many also incorporate masks to specify the editing region (Wang et al., 2023; Zhao et al., 2024). However, due to the inherent imprecision of both prompts and masks, some methods integrate additional inputs to provide finer control over the generation process (Yang et al., 2024; Shi et al., 2024b; Liu et al., 2024a).
Among these, two methods are most closely related to our workflow. First, MagicQuill (Liu et al., 2024b) enables user-guided editing using color and edge information. However, its color guidance is constrained to low-resolution 32$\times$32 blocks, regardless of image size, preventing fine-grained edits such as those required in EditEval. More discussion and a case study to compare SPICE and MagicQuill is included in Appendix J. Secondly, self-guidance (Epstein et al., 2023) allows users to specify object placement and size through textual prompts, but this approach remains inherently imprecise due to the limitations of language-based descriptions. SPICE overcomes the constraints of the two methods by accepting full-resolution color and edge hints and synergizing two complementary diffusion models. Hence, SPICE effectively interprets and applies the provided guidance, leading to precise and high-quality edits.
Image Editing Benchmarks
Several datasets have been introduced for training and benchmarking image editing models, including EditBench (Wang et al., 2023), MagicBrush (Zhang et al., 2024), HQ-Edit (Hui et al., 2024), InstructPix2Pix (Brooks et al., 2023), UltraEdit (Zhao et al., 2024), Seed-Data-Edit (Ge et al., 2024), and EditEval (Huang et al., 2024). While dataset sizes have grown over time, the number of challenging test cases remains limited. Future benchmarks should prioritize structural editing tasks, such as modifying object actions or layouts, rather than focusing primarily on simpler semantic edits like color or texture changes. Simple semantic edits can already be handled almost perfectly by our workflow.
5 Conclusion
Existing prompt-based image editing models fail in performing local edits, working under different resolutions, following user instructions, and maintaining image quality during multiple editing steps. We propose SPICE, a training-free workflow that addressees all challenges above. We release the workflow to facilitate future research.
Impact Statement
For the first time, our workflow unlocks the ability to perform photo-realistic edits with perfect details using diffusion models. Since the workflow is fully open source, malicious users may misuse the workflow, leading to negative societal consequences. However, as there exist various techniques for detecting AI-generated content, we do not believe the edited images will pose an immediate threat to the public.
Appendix A Baseline Model Hyperparameters
@@
For InstructPix2Pix, we use the model and hyperparameters according the Hugging Face official page [^14] https://huggingface.co/timbrooks/instruct-pix2pix. The number of inference steps is set to 10, and the image guidance scale is set to 1.
@@
For MagicBrush, we also refer to the Hugging Face page [^15] https://huggingface.co/osunlp/InstructPix2Pix-MagicBrush. Specifically, we use the recommended “recent best checkpoint”, which is MagicBrush-epoch-52-step-4999.ckpt. We do not provide extra hyperparameters when calling the inference script from the command line, except that we provide a random seed. Hence, all parameters are fixed at their default values, presumably recommended by the model developers.
@@
For UltraEdit, we first use the model version that supports masks [^16] https://huggingface.co/BleachNick/SD3_UltraEdit_w_mask. We use the recommended hyperparameter values. We also use the free-form version [^17] https://huggingface.co/BleachNick/SD3_UltraEdit_freeform, since binary masks are not available in Emu Edit and MagicBrush test sets. Since there is not a recommended set of hyperparameters on the official page for the free-form version, we use the same hyperparameters for the free-form version and the masked version.
Appendix B Human Evaluation Details

### v2

@@ B.1 Benchmark Results
We use InstructPix2Pix (IP2P) [3], IP2P trained on MagicBrush [29], and UltraEdit [32] as baseline methods. We evaluate all models on the EditEval [10] benchmark, using the standard CLIP [19] text-image direction similarity (CLIPdir) and CLIP output similarity (CLIPout) metrics from the Emu Edit benchmark [20]. More details can be found in Appendix˜C.
@@
Table˜2 shows that our workflow achieves the highest scores on the EditEval benchmark. In addition, we compare the edited images by baseline methods and our workflow on all six different editing categories for visual comparison. Figure˜5 shows the failure patterns of baseline methods including undesired global color shift of the grass (Object Addition), removing the lens flare outside the masked region (Object Replacement), inability to remove multiple humans from the image (Object Removal), and wrong anatomy of the polar bear (Action Change). In contrast, our workflow maintains global color, keeps details outside the mask untouched, removes objects as requested, and generates animals with correct anatomy.
Table 2: Our workflow achieves top performance in two quantitative metrics on the EditEval benchmark. We show mean $\pm$ standard deviation calculated across all images ($n=126$).
@@
To provide a comprehensive qualitative evaluation, we further conduct a human evaluation to compare results from different models (more details in Appendix˜D). Figure˜6 shows that our workflow is predominantly preferred. Although at a much lower chance than baseline methods, our workflow does sometimes fail to follow instructions, as indicated by the “both bad” cases. In Section˜B.3, we show how the flexibility of our workflow allows us to overcome this limitation easily.
Figure 6: Our workflow is preferred by the annotators over any baseline method. In this evaluation, we use only a single editing step and fixed hyperparameters for our workflow, and thus our workflow can still fail (“both bad” cases). In Section˜B.3, we show how relaxing these constraints can dramatically improve results.
@@
In Appendix˜E, we show that our workflow outperforms baselines on challenging examples from two other popular benchmarks (Emu Edit and MagicBrush). On the challenging examples, our workflow also outperforms GPT-4o, Gemini 2.0 Flash, and SeedEdit [21] (a recent mask-based commercial baseline from Doubao AI).
@@
While FLUX.1 [dev] is a strong image generation backbone, we demonstrate that our workflow outperforms the backbone model FLUX.1 [dev] by an ablation study. Also, we designed our workflow such that it mitigates many issues of the specialized inpainting model FLUX.1 [dev] Fill. [^11] https://huggingface.co/black-forest-labs/FLUX.1-Fill-dev Hence, the better performance of our workflow than baseline methods does not result from simply selecting a stronger backbone. Results can be found in Appendix˜F.
@@
Despite the superior performance, our workflow imposes minimal burden on the users. While the users need to provide both the masks and the hints, the two inputs can be casually sketched, because our workflow is robust to missing details and inaccurate shapes. Examples of excellent editing results from simple user inputs are shown in Appendix˜G.
Image Editing Benchmarks.
Several datasets have been introduced for training and benchmarking image editing models, including EditBench [25], MagicBrush [29], HQ-Edit [11], InstructPix2Pix [3], UltraEdit [32], Seed-Data-Edit [8], and EditEval [10]. While dataset sizes have grown over time, the number of challenging test cases remains limited. Future benchmarks should prioritize difficult structural editing tasks, such as modifying object actions or layouts, rather than focusing primarily on simpler semantic edits like color or texture changes.
@@
Figure 7: Our workflow delivers precise and repeatable edits, where the generated objects align with user specifications. The top images compare the user-specified color hints with the generated objects, and the bottom plots show the percentage errors. On each plot, we show the mean and standard deviation of percentage errors across 10 random seeds. All mean errors are close to 0.
B.2 Precise Editing
@@
To quantify the precision of our workflow, we add objects while specifying 5 properties, including size, location, rotation, color, and aspect ratio (Figure˜7). To specify the value of each property, we use a hint in the form of a color patch. In the edited image, we measure the property value of the added object and calculate the percentage error against the user-specified property value (Appendix˜H). We use 10 random seeds to account for generative variability.
@@
Figure˜7 shows that the mean of percentage errors for each specified property value is close to 0, with a small standard deviation across random seeds. There are still systematic errors, such as the height of the apple being consistently 10% larger than specified. However, for the apple, the height is larger because the apple’s stem is unspecified by the hint. Overall, the results confirm that our workflow delivers precise, repeatable edits critical for real-world scenarios.
B.3 Iterative Editing
@@
In this subsection, we consider two iterative editing regimes, one using fewer than 5 editing steps, another more than 100. Multiple editing steps are often necessary, as one single editing or generation step usually fails to align with complicated user requirements (further discussed in Section˜B.4).
@@
In the low-steps regime, if unsatisfied with results from the first editing step, we can improve them using another editing step with our workflow. In Section˜B.1, we identify that our workflow does not work well for a few cases in single-step editing. In these cases, a large proportion of the edited region is satisfactory, leaving only a small proportion to be fixed. For baseline methods, as the image quality degrades after the first editing step, we cannot run a second editing step but can only rerun the model with the same input and different random seeds, and the results are unsatisfactory (Appendix˜I). However, with our workflow, we can improve the result by recycling the first output (Figure˜8).
@@
In the high-steps regime, while many previous methods support multi-step editing, heavy artifacts are generated and propagated in each step, disqualifying these methods for long editing tasks. For example, when GPT-4o is prompted to return the same image to the user, [^12] https://replicateimage.com/examples the image quality quickly deteriorates within even a few steps, disqualifying GPT-4o from iterative editing. In contrast, our workflow steadily improves the image after multiple steps. In Figure˜9(a), we show a result after 40 diverse editing steps, including outpainting, structural edits, upscaling, relighting, and composition adjustment (forming a heart shape using the sky, the crescent, the claw, and the branches). Full steps are shown in Appendix˜J. Interestingly, to relight characters according to the environment, we do not need to provide any color hints. We also do not need a specialized relighting model, such as IC-Light [31]. In Figure˜9(b), we start with a model-generated image with various errors in shadows, objects, and anatomy. Using the same model as the backbone in our workflow, we fix these errors one by one across more than 100 editing steps. This image is produced by BoleroMix (Pony) v1.41, a checkpoint derived from SDXL. More examples of final edited results in different styles can be found in Appendix˜K, where around 200 editing steps are performed to achieve the desired level of detail. Each image is created from scratch within 4 hours in one sitting.
Original Step 1 Step 2 Step 3 Step 4
@@
Step 1 Step 40
Refer to caption Refer to caption
(a) Iteratively constructing an image
Step 1 Step 100+
Refer to caption Refer to caption
(b) Iteratively refining an image
Figure 9: We use SPICE to consistently improve image quality during a large number of editing steps. The editing steps can either change the structure or refine the details. The top images are generated using Flux.1 [dev]. The bottom images are generated using BoleroMix (Pony) v1.41.
B.4 Customizable Editing
@@
Prompt-based image generation or editing models frequently fail to fulfill their promise that users can create content according to their own will. Users are frustrated for two main reasons.
@@
First, prompts are hard to design and often cannot instruct the model to produce complex output. A user assumes that the model follows the prompt, but the model does not interpret the prompt as humans do. One example is that DALL·E 3 [1] in November 2024 still fails to generate a rabbit with four ears or a violin without a bridge, [^13] https://cs.nyu.edu/~davise/papers/DALL-E-Parts/PartsNovember24/DALL-E-Parts-November24.html as shown in Figure˜10(a). Another example is that DALL·E 3 fails to interpret relations or numbers, such as a potato under a spoon or five fish [4].
Round 1 Round 1
Refer to caption Refer to caption
Round 2 Round 2
Refer to caption Refer to caption
Round 3 Round 3
Refer to caption Refer to caption
(a) DALL·E 3
(b) GPT-4o
(c) Ours
Figure 10: SPICE can generate content that DALL·E 3 and GPT-4o cannot. Two examples are a rabbit with 4 ears and a violin without a bridge. By January 2025, DALL·E 3 fails even after the user points out the error and asks the model to edit the errors multiple times. By May 2025, GPT-4o also cannot generate all elements correctly at the same time. In contrast, our workflow generates all elements correctly in one image, demonstrating superior customizability.
@@
Second, the innate randomness of the model dramatically increases the generation cost or even prohibits the generation of complicated images. Suppose a user wants to generate an image with 10 different objects at various locations on this image. If each object has a 50% chance to be generated perfectly, the success rate of all objects being perfect is below 0.1%. In real scenarios, the perfection rate for each object will only be lower, and the number of objects larger. Then, generating such an image is practically impossible.
With SPICE, instead of desperately engineering prompts or varying the random seed, users handle these difficulties by tuning 3 hyperparameters, namely context size, denoising strength, and Canny model steps. By slightly tuning the 3 hyperparameters (Appendix˜L), we are able to generate Figure˜10(c), an image with a rabbit of four ears, holding a violin without a bridge with the left hand and a spoon with the right hand. There is a potato under the spoon and five fish in the air. [^14] Figure 10(c) has been accepted to CVPR AI Art Gallery 2025. https://thecvf-art.com/project/compositionality-and-parts/ In contrast, DALL·E 3 and GPT-4o both fail to generate the components correctly. The prompt for GPT-4o is “Generate a rabbit with four ears holding a violin without a bridge. The rabbit holds a spoon in her hand. There is a potato under the spoon. There are five fish floating in the air.”
Appendix C Evaluation Details
Evaluation Benchmark.
We use the second version of the EditEval [10] benchmark to evaluate our workflow and baseline methods. EditEval is a single-step editing benchmark consisting of original images, source captions, target captions, and editing prompts. The original images have resolutions ranging from $\times1863$ to $\times8742$. EditEval covers semantic editing (object addition, object removal, object replacement, and background change), stylistic editing (style change and texture change), and structural editing tasks (action change). Due to the extremely high resolutions of images, many editing tasks in EditEval require challenging fine-grained edits. An example is removing a tiny insect from a bird’s beak, which requires editing fewer than 1% of all pixels. EditEval allows a fair comparison of all methods, as none of them were trained on a dataset with the same distribution as EditEval.
Baseline Methods.
The baseline methods include InstructPix2Pix (IP2P) [3], IP2P trained on MagicBrush [29], and UltraEdit [32]. For these methods, we use the original editing prompts from EditEval. The inference hyperparameters are set to recommended values. For UltraEdit, we use the mask-based checkpoint and the same binary masks (with context dots) as in our workflow. Hence, UltraEdit similarly benefits from the extra user input, ensuring a fair comparison between our workflow and this strong baseline.
@@ Appendix C Evaluation Details
For InstructPix2Pix, we use the model and hyperparameters according the Hugging Face official page. [^15] https://huggingface.co/timbrooks/instruct-pix2pix The number of inference steps is set to 10, and the image guidance scale is set to 1.
@@
For MagicBrush, we also refer to the Hugging Face page. [^16] https://huggingface.co/osunlp/InstructPix2Pix-MagicBrush Specifically, we use the recommended “recent best checkpoint”, which is MagicBrush-epoch-52-step-4999.ckpt. We do not provide extra hyperparameters when calling the inference script from the command line, except that we provide a random seed. Hence, all parameters are fixed at their default values, presumably recommended by the model developers.
@@
For UltraEdit, we first use the model version that supports masks. [^17] https://huggingface.co/BleachNick/SD3_UltraEdit_w_mask We use the recommended hyperparameter values. We also use the free-form version, [^18] https://huggingface.co/BleachNick/SD3_UltraEdit_freeform since binary masks are not available in Emu Edit and MagicBrush test sets. Since there is not a recommended set of hyperparameters on the official page for the free-form version, we use the same hyperparameters for the free-form version and the masked version.
Evaluation Metrics.
We use the CLIP [19] text-image direction similarity (CLIPdir) and CLIP output similarity (CLIPout) metrics from the Emu Edit benchmark [20]. These two metrics are suitable for reference-image-free evaluation. We do not calculate the L1 distance between edited and original images, as the L1 distance does not monotonically increase with the editing quality (higher distance can simultaneously indicate better global editing performance and worse local editing performance). In EditEval, we exclude the style change task, as the task is better handled by style LoRAs or specialized checkpoints (e.g., Flux.1 [dev] Redux). [^20] https://huggingface.co/black-forest-labs/FLUX.1-Redux-dev After excluding 24 images, there are $n=126$ images remaining.
Appendix D Human Evaluation Details